# Large Language Model Can Continue Evolving From Mistakes

### v1

@@ 4.1 Experimental Settings
Base language models. The objective of our work is to evaluate the effectiveness of the CEM method in enhancing the performance of LLMs, even those with smaller parameter scales. We select open-source models with parameter ranges from 6 to 13 billion for our experiment. We use CuteGPT-13B [^4] https://huggingface.co/XuYipei/kw-cutegpt-13b-base, ChatGLM-6B [^5] https://huggingface.co/THUDM/chatglm3-6b, Qwen-7B [^6] https://huggingface.co/Qwen/Qwen-7B-Chat as our instruction fine-tuning base models for CEM. These models have been fine-tuned on the instruction data enabling them to effectively follow instructions. CuteGPT is based on the original Llama model structure, expands the Chinese vocabulary and performs pre-trainingLiang et al. (2023). ChatGLM-6B also is an open bilingual language model. It is trained for about 1T tokens of Chinese and English corpusDu et al. (2022). Qwen-7B-Chat is a Transformer-based large language model, which is pretrained on a large volume of data, including web texts, books, codes, etcBai et al. (2023).
In the instruction fine-tuning stage, CuteGPT-13B and Qwen-7B-chat employ full fine-tuning, and ChatGLM-6B employs Prompt-tuningLiu et al. (2022).
Supplementary training dataset. We construct a variety of supplementary training sets and their respective control groups for the experiments following the method described in Section 3.3. To distinguish between the model using different supplementary training sets, we establish the following naming rules:
(1) Baseline: it represents the model that has not undergone supplementary training and has only been pre-finetuned.
(2) N: N denotes Normal, indicating that the model is supplementary trained using the supplementary training set constructed by the Normal Method.
(3) E and EC: E denotes Extract, indicating that the model is supplementary trained using the supplementary training set constructed by the Extract Method. To obtain its respective control group, Extract Control (EC), we replace the Extract Corpus in the training set with the General Corpus.
(4) R1, R2, R3 and RC: R denotes Review, indicating that the model is supplementary trained using the supplementary training set constructed by the Review Method and based on the proportion of ’Format Corpus’ and ’Extract Corpus’ formats in the Review Corpus, we further name them as R1, R2, and R3. To obtain their respective control group, Review Control (RC), we replace the Review Corpus in the training set with the General Corpus.
In addition, we distinguish between the data sources from which the supplementary corpus is collected and establish the following naming rules:
(1) Wiki: it represents the Supplementary Corpus of the supplementary training set used by the model is from Wikipedia.
(2) Bing: it represents the Supplementary Corpus of the supplementary training set is from Bing.
(3) Mix: it represents the Supplementary Corpus of the supplementary training set is collected from both Bing and Wikipedia and concatenated.
@@
Metrics. The effects of supplementary training will be evaluated using the following metrics:
(1) ACC: it represents the accuracy of the answer of LLM.
@@
(2) W2R: it indicates the percentage of questions where the LLM changes incorrect answers to correct ones in the test set.
@@
(3) R2W: it indicates the percentage of questions where the LLM changes correct answers to incorrect ones in the test set.
In Section 4.2, we combine the above two naming rules to describe the experimental results clearly. For instance, Wiki_E denotes supplementary training using the dataset constructed through the Extract Method. The Supplementary Corpus of this dataset is collected from Wikipedia.
4.2 Results and Analysis
@@
Figure 3: This figure shows the accuracy of various models using the CEM method on the Xiezhi and CMMLU tasks. The supplementary training set is constructed using the Normal Method.
CEM method enhances the performance of various LLMs across diverse tasks. We conduct tests on LLMs that undergo supplementary training using the CEM method. The training sets are constructed via the Normal Method. The accuracy results are presented in Fig. 3. More complete experimental results are in Appendix Table 4 and 5. After implementing CEM, all LLMs showed significant improvement in accuracy for both QA tasks, demonstrating the effectiveness of CEM in addressing knowledge deficiencies and enhancing LLMs’ performance. For instance, after one round of of CEM optimization with Bing as its supplementary corpus source, the accuracy of Qwen-7B-Chat on the CMMLU task surged by 16.08%.
@@
Smaller LLMs can better benefit from CEM. The effectiveness of LLM tuning with CEM depends on the innate abilities of the LLM and also varies for LLMs with different parameter scales. According to the experimental results presented in Fig. 3, among the open-source models mentioned, ChatGLM-6B, which has the fewest parameters, shows the most significant improvement, up to 13.17% on the Xiezhi task. In contrast, CuteGPT-13B, which has the highest initial answer accuracy, exhibits a smaller improvement of up to 3.63% compared to the smaller LLM.
Extract Corpora and Review Corpora can significantly boost accuracy. We conducted parallel experiments on the supplementary training set construction method described in section 3.3, using CuteGPT-13B on Xiezhi dataset. The results are presented in Fig. 4. CEM with Extract Corpus outperforms its control group in all cases with higher Accuracy and W2R metric, e.g., Mix_E shows a 3.11% increase in accuracy to its control group Mix_EC, and a 10.96% increase compared to the baseline.
@@
Figure 4: The figure shows the experimental results of CuteGPT-13B on the Xiezhi task after being trained with a supplementary training set constructed using various methods. ’Avg of Ri’ and ’Best of Ri’ represent the average and best results, respectively, of the three Review Methods.
Meanwhile, in alignment with expectations, CEM with the Review Corpus ($C_{R_{i=1,2,3}}$), consistently outperform their control group across all instances and significantly reduce the R2W metric. For example, Wiki_R1 shows a 5.74% improvement in accuracy and a 3.70% reduction in R2W compared to its control group, Wiki_RC, and a 12.25% improvement in accuracy and a 5.89% reduction in R2W compared to the baseline. The results indicate that the W2R is improved by the Extract Corpus. This suggests that the model’s capacity to identify and comprehend the corresponding knowledge in the corpus is enhanced, leading to better error correction. Review Corpus significantly reduced R2W, indicating that the model’s forgetting of correct knowledge was suppressed.
In addition, the three construction methods, $R_{1}$,$R_{2}$ and $R_{3}$, are similar in their ability to enhance LLMs’ performance, which means either form of Review Corpus can be effective. Therefore, if not specified, $R$ will refer to the $R_{1}$ Method in the following.
@@
Supplementary corpora from multiple sources do not always improve the CEM approach more than a single source. The effectiveness of CEM depends on the recall rate when searching corpora and the quality of supplementary corpora, as this determines whether the corpus contains knowledge that the LLM lacks. Moreover, noise and errors within low-quality corpora can negatively impact the LLM. Results in Fig. 4 show that there is no significant difference in the effectiveness of CEM, whether using Wikipedia or Bing as a single data source, or when using multiple data source concatenated. When using the supplementary training set constructed by Review Method, Mix_R has lower accuracy than Wiki_R and Bing_R despite having more data volumn, This may be due to the information redundancy caused by the high overlap of the supplementary corpus from different sources. Additionally, an excess of supplementary corpus may dilute the proportion of generalization, excerpt, and comment corpus, thereby hindering their effectiveness in reducing model catastrophic forgetting and enhancing model knowledge extraction.
@@
W2R comp. to Baseline 24.46 26.03 28.29 22.07 23.32 27.01
comp. to Round1 20.00 17.56 16.43 18.49 18.33 16.61
R2W comp. to Baseline 15.70 13.76 11.29 18.63 17.51 11.61
comp. to Round1 14.87 14.39 11.68 18.03 18.06 12.14
Table 3: We employ the identical training data construction method as in the first round to conduct a second round of supplementary training on the LLM. The table shows the experimental results of CuteGPT-13B using the CEM method on the Xiezhi task.
CEM can continuously evolve the LLM through multiple iterations. We conduct a second round of supplementary training on CuteGPT-13B on Xiezhi task using the same dataset construction method and data source as the first round. The results are presented in Table 3. After the second round of supplementary training, the accuracy of LLM’s answer on both two QA tasks has significantly improved compared to the previous round. Both the Extract Method and Review Method are effective in improving the accuracy of the model. For instance, Wiki_R shows a 4.75% improvement in accuracy compared to its previous round, and a 17.00% improvement in accuracy compared to baseline.
Furthermore, the Review Method consistently exhibits significantly lower R2W metric compared to other supplementary dataset construction methods, e.g., the R2W metric relative to Baseline of Wiki_R decreases by 7.02% compared to Wiki_N and 5.90% compared to Wiki_E. We argue that the Review Method performs best because it significantly reduces the probability of changing a previously correct response to an incorrect one. This enables the model to be evolved by two or more rounds of supplemental training.
@@ 5 Conclusion
In this paper, we propose and validate the Continue Evolving from Mistakes (CEM) method, which assesses LLM’s knowledge deficiencies through LLM’s mistakes and targets supplementary LLM’s knowledge. We first devise a general process for supplementary training: (1) testing LLM and collecting mistakes, (2) collecting supplementary corpora from multiple data sources, (3) constructing a supplementary training set and fine-tuning the LLM. Furthermore, we introduce multiple methods for constructing training data to enhance the LLM’s capability to extract and comprehend knowledge within the corpora, and to mitigate the forgetting of LLM. We conduct experiments on several opensource LLMs to validate the efficacy of CEM. The results consistently show that CEM supplements the LLM’s knowledge deficiencies and improves the accuracy of the answers. For instance, on the Xiezhi task, the accuracy of the CuteGPT-13B improves by up to 12.25% in the first round and up to 17.00% in the second round.
Limitations: This paper focuses on the model’s performance on question-answering task. It is yet to be determined if CEM can consistently enhance the model’s other abilities and its performance in non-QA tasks.
In future research, we will further explore three or more rounds of supplementary training to improve the LLM and to mitigate the catastrophic forgetting. In addition, we will experiment with other data sources for corpora and optimize supplementary corpus collection strategies to equip the LLM with more advanced continual iterative evolutionary capabilities.
@@ Appendix A Process of Collecting Corpora from Multiple Sources
Fig.5 and Fig.6 show the process of obtaining supplementary data from Wikipedia and Bing respectively.
@@ Appendix B Experimental Results of Various LLMs on QA tasks
Table 4 and Table 5 show the complete experimental results of LLMs using the CEM method on Xiezhi and CMMLU tasks. We use the Normal method to construct the supplementary training dataset.
@@
Baseline Wiki Bing Baseline Wiki Bing Baseline Wiki Bing
@@
Baseline Wiki Bing Baseline Wiki Bing Baseline Wiki Bing
@@
Appendix C Experimental Results with Various Supplementary Training Dataset Construction Methods
Table 6 shows the complete experimental results of CuteGPT-13B on the Xiezhi task after being trained with a supplementary training set constructed using various methods.
Data Source Metrics Baseline N E EC R1 R2 R3 RC
@@ Appendix C Experimental Results with Various Supplementary Training Dataset Construction Methods
Table 6: The table shows the experimental results of CuteGPT-13B on the Xiezhi task using different supplementary training set construction methods.
@@ Appendix D Data Division
Table 7 shows the the data division of Xiezhi and CMMLU. The multiple-choice questions are divided into three parts. The Pre-Finetuning part is utilized to train the model to answer questions in a standard format before supplementary training. The Testing part is used to construct the test set, and the remaining part is used to construct the Format Corpus and Extract Corpus for supplementary training.
Dataset Total Pre-Finetuning Testing Supplementary Training

### v3

@@ 4.2 Base Language Models
To evaluate the effectiveness of the CEM method in enhancing the performance of LLMs, even for those with smaller parameter scales, we select open-source LLMs with parameter ranges from 6 to 13 billion. Our instruction fine-tuning base models include CuteGPT-13B [^4] https://huggingface.co/XuYipei/kw-cutegpt-13b-base, ChatGLM-6B [^5] https://huggingface.co/THUDM/chatglm3-6b, and Qwen-7B-Chat [^6] https://huggingface.co/Qwen/Qwen-7B-Chat. These models have undergone fine-tuning on the IFT data to ensure they effectively follow instructions.
CuteGPT-13B is based on the original Llama model structure, expands the Chinese vocabulary and performs pre-trainingLiang et al. (2023).
ChatGLM-6B is an open bilingual language model. It is trained for about 1T tokens of Chinese and English corpusDu et al. (2022).
Qwen-7B-Chat is a Transformer-based large language model, which is pretrained on a large volume of data, including web texts, books, codes, etcBai et al. (2023).
4.3 Metrics
@@
The effectiveness of supplemental training will be assessed using the following metrics:
ACC: it represents the accuracy of the answer of LLM.
@@
W2R: it indicates the percentage of questions where the LLM changes incorrect answers to correct ones on the test set.
@@
R2W: it indicates the percentage of questions where the LLM changes correct answers to incorrect ones on the test set.
5 Experiments
We construct various supplemental training sets and their corresponding control groups for the experiments, following the method described in Section 3.3. To differentiate between models trained on different supplemental training sets, we establish the following naming conventions:
INIT: it represents the model, which has undergone only pre-finetuning using the Normative Corpus, without any supplemental training.
CEM-N: N denotes Normative, representing the model that has undergone supplemental training using the training set constructed by the Plain Strategy.
[…]
We append suffixes to the aforementioned naming conventions to differentiate the data sources of the supplemental corpus. These suffixes include: ‘Wiki’, ‘Bing’ and ‘Mix’ where ‘Mix’ indicates that the Supplemental Corpus is collected from both Bing and Wikipedia and concatenated. For instance, ‘E_Wiki’ denotes the model supplemented trained using the dataset constructed through the Extractive Strategy. The Supplemental Corpus of this dataset is collected from Wikipedia.
5.1 Main Results
We conduct experiments to analyze the accuracy after applying CEM on the test data. The results are shown in Figure 2. Our observations are summarized as follows:
@@ 5.1 Main Results
Figure 2: This figure shows the accuracy of various models using the CEM method on the Xiezhi and CMMLU tasks. The supplemental training set is constructed using the Plain Strategy.
Firstly, following the supplemental training, a noticeable enhancement in accuracy is observed across all models for the two given tasks. This implies that despite the high availability of supplemental data and the small size of the training set, the CEM method effectively complements the model’s knowledge deficiencies. For instance, the ChatGLM-6B model shows an improvement of up to 16.08% on the CMMLU task.
@@
Secondly, the effectiveness of tuning LLMs with CEM depends on their innate abilities and also varies across LLMs with different parameter scales. Among the mentioned open-source models, ChatGLM-6B, which has the fewest parameters, exhibits the most significant improvement, up to 13.17% on the Xiezhi task. In contrast, CuteGPT-13B, with the highest initial answer accuracy and the most parameters, shows a smaller improvement of up to 3.63% compared to the smaller LLM.
Construction Strategy Analysis. To investigate the effectiveness of each component in supplemental training set, we conduct experiments on LLMs trained using various dataset construction strategies, the results are shown in Figure 3. Our observations are summarized as follows:
@@
Figure 3: The figure shows the experimental results of CuteGPT-13B on the Xiezhi task after being trained with a supplemental training set constructed using various strategies. ’Avg of $R_{\alpha}$’ and ’Best of $R_{\alpha}$’ represent the average and best results, respectively, of the Review Strategy with different $\alpha$.
Firstly, Extractive Corpus and Review Corpus can significantly boost LLMs’ performance. CEM with Extractive Corpus ($C_{E}$) outperforms its control group in all cases with higher Accuracy and W2R metric, e.g., E_Mix shows a 3.11% increase in accuracy to its control group E’_Mix, and a 10.96% increase compared to INIT. Meanwhile, in alignment with expectations, CEM with the Review Corpus ($C_{R_{\alpha=0,0.5,1}}$), consistently outperform their control group across all instances and significantly reduce the R2W metric. For example, R0.5_Wiki shows a 5.74% improvement in accuracy and a 3.70% reduction in R2W compared to its control group, R’_Wiki, and a 12.25% improvement in accuracy and a 5.89% reduction in R2W compared to the INIT.
The results indicate an improvement in W2R metric due to the use of the Extractive Corpus, suggesting an enhanced capacity of the model to identify and comprehend relevant knowledge within the corpus. This enhancement leads to more effective error correction. Additionally, the Review Corpus significantly reduced R2W, indicating suppression of the model’s tendency to forget correct knowledge.
@@
Secondly, Supplemental Corpus from multiple sources do not always improve the CEM method more than a single source. The effectiveness of CEM depends on the recall rate when searching corpus and the quality of supplemental corpus, as this determines whether the corpus contains knowledge that the LLM lacks. Moreover, noise and errors within low-quality corpora can negatively impact the LLM. Results in Figure 3 show that there is no significant difference in the effectiveness of CEM, whether using Wikipedia or Bing as a single data source, or when using multiple data source concatenated. When using the supplemental training set constructed by Review Strategy, R_Mix has lower accuracy than R_Wiki and R_Bing despite having more data volumn, This may be due to the information redundancy caused by the high overlap of the supplemental corpus from different sources. Additionally, an excess of supplemental corpus may dilute the proportion of General, Extractive, and Normative Corpus, thereby hindering their effectiveness in reducing model catastrophic forgetting and enhancing model knowledge extraction.
Thirdly, proportion factor $\alpha$ has limited impact on the effectiveness of the Review Corpus. According to Table 6 in Appendix, R0, R0.5 and R1 are similar in their ability to enhance LLMs’ performance. This may be because the models themselves have already mastered the knowledge, and replaying it in either format serves to consolidate that knowledge. Therefore, if not specified, R will refer to the R0.5 Strategy in the following.
5.2 Multiple Iterations Analysis
@@
W2R vs. INIT 24.46 26.03 28.29 22.07 23.32 27.01
vs. Round1 20.00 17.56 16.43 18.49 18.33 16.61
R2W vs. INIT 15.70 13.76 11.29 18.63 17.51 11.61
vs. Round1 14.87 14.39 11.68 18.03 18.06 12.14
Table 3: We employ the identical training data construction strategy as in the first round to conduct a second round of supplementary training on the LLM. The table shows the experimental results of CuteGPT-13B using the CEM method on the Xiezhi task.
We conduct a second round of supplemental training on CuteGPT-13B on Xiezhi task using the same dataset construction strategy and data source as the first round. The results are presented in Table 3. After the second round of supplementary training, the accuracy of LLM’s answer on both two QA tasks has significantly improved compared to the previous round. Both the Extractive Strategy and Review Strategy are effective in improving the accuracy of the model. For instance, R_Wiki shows a 4.75% improvement in accuracy compared to its previous round, and a 17.00% improvement in accuracy compared to INIT.
Furthermore, the Review Strategy consistently exhibits significantly lower R2W metric compared to other supplementary dataset construction strategies, e.g., the R2W metric relative to INIT of Wiki_R decreases by 7.02% compared to Wiki_N and 5.90% compared to Wiki_E. We argue that the Review Strategy performs best because it significantly reduces the probability of changing a previously correct response to an incorrect one. This enables the model to be evolved by two or more rounds of supplemental training.
6 Discussion
Why can the Extractive Corpus be effective? We observe that models employing the Extractive Corpus exhibit higher accuracy and W2R metric. Unlike the Review Corpus, where improving model performance is intuitive, the precise mechanism by which the Extractive Corpus augments the model’s capability to capture pivotal knowledge in lengthy texts remains worthy of discussion.
Extractive Corpus provides the model with both questions and supplemental material containing relevant knowledge. Questions and options typically present straightforwardly and can be condensed into several key points, whereas the supplemental material tends to intricately interweave numerous factual statements Jiang et al. (2024). Thus, we suggest that questions may be a structured induction of supplemental material, where the Extractive Corpus explicitly links complex information within the corpus to the resulting induction (i.e., the question and options). Questions and options can act as cues and guides for the model to gather relevant information from the corpus, considering how this knowledge is accessed through questions. This, in turn, can enhance the model’s ability to extract useful information from the corpus while eliminating redundant and irrelevant information.
@@ 7 Conclusion
In this paper, we propose and validate the Continue Evolving from Mistakes (CEM) method, which assesses LLM’s knowledge deficiencies through LLM’s mistakes and targets supplementary LLM’s knowledge. We first devise a general process for supplementary training: (1) testing LLM and collecting mistakes, (2) collecting supplemental corpora from multiple data sources, (3) constructing a supplementl training set and fine-tuning the LLM. Furthermore, we introduce multiple strategies for constructing training data to enhance the LLM’s capability to extract and comprehend knowledge within the corpora, and to mitigate the forgetting of LLM. We conduct experiments on several open-source LLMs to validate the efficacy of CEM. The results consistently show that CEM supplements the LLM’s knowledge deficiencies and improves the accuracy of the answers. For instance, on the Xiezhi task, the accuracy of the CuteGPT-13B improves by up to 12.25% in the first round and up to 17.00% in the second round.
8 Limitations
This paper focuses on the model’s performance on question-answering task. It is yet to be determined if CEM can consistently enhance the model’s other abilities and its performance in non-QA tasks.
In future research, we will further explore three or more rounds of supplemental training to improve the LLM and to mitigate the catastrophic forgetting. In addition, we will experiment with other data sources for corpora and optimize supplementary corpus collection strategies to equip the LLM with more advanced continual iterative evolutionary capabilities.
@@ Appendix A Process of Collecting Corpora from Multiple Sources
Fig.4 and Fig.5 show the process of obtaining supplemental data from Wikipedia and Bing respectively.
@@ Appendix B Experimental Results of Various LLMs on QA tasks
Table 4 and Table 5 show the complete experimental results of LLMs using the CEM method on Xiezhi and CMMLU tasks. We use the Plain Strategy to construct the supplemental training dataset.
@@
INIT Wiki Bing INIT Wiki Bing INIT Wiki Bing
@@
INIT Wiki Bing INIT Wiki Bing INIT Wiki Bing
@@
Appendix C Experimental Results with Various supplemental Training Dataset Construction Strategies
Table 6 shows the complete experimental results of CuteGPT-13B on the Xiezhi task after being trained with a supplemental training set constructed using various strategies.
Data Metrics INIT CEM-N CEM-E CEM-E’ CEM-R0.5 CEM-R1 CEM-R0 CEM-R’
@@ Appendix C Experimental Results with Various supplemental Training Dataset Construction Strategies
Table 6: The table shows the experimental results of CuteGPT-13B on the Xiezhi task using different supplemental training set construction methods.
@@ Appendix D Data Division
Table 7 shows the the data division of Xiezhi and CMMLU. The multiple-choice questions are divided into three parts. The Pre-Finetuning part is utilized to train the model to answer questions in a standard format before supplemental training. The Testing part is used to construct the test set, and the remaining part is used to construct the Normative Corpus and Extractive Corpus for supplemental training.
Dataset Total Pre-Finetuning Testing supplemental Training